OpenStudy (anonymous):

The sum of three times a number and a second number will total 12 or greater. The sum of three times the first number and twice the second number is less than or equal to 15. Find the two numbers. A. first number = 2 and second number = -1 B. first number = 3 and second number = 2 C. first number = 5 and second number = 1 D. first number = 4 and second number = 1

OpenStudy (raden):

the are 2 systems : 3x +y >= 12 3x + 2y <= 15 trial errors is works, just identify all options. which points be right ?

OpenStudy (raden):

for example, on option A, point of (x,y) = (2,-1) put point (2,-1) into the sytems : 3x +y >= 12 and 3x + 2y <= 15 check it, 3(2) + (-1) >= 12 6 - 1 > = 12 5 >= 12 it is FALSE so, we neednt check in the 2nd system, because in the 1st system already be false.

OpenStudy (anonymous):

It's c.

OpenStudy (raden):

let's check together if it is C (5,1) 3(5) +1 >= 12 15+1>= 12 16>= 12 RIGHT 3(5) + 2(1) <= 1 15 + 2 <= 1 17 <= 1 it is FALSE try again :)

OpenStudy (anonymous):

B?

OpenStudy (raden):

no, it is D (4,1) 3(4) +1 >= 12 12+1>= 12 13>= 12 it is RIGHT 3(4) + 2(1) <= 15 12 + 2 <= 15 14 <= 15 it is RIGHT so, B is the answer

OpenStudy (raden):

opsss .. i mean D is the answer

OpenStudy (anonymous):

Dang I already had that I am retesting haha. Thank you so much.

OpenStudy (raden):

you're welcome :*